OpenStudy (anonymous):

A bag contains 30 red tiles, 15 green tiles, and 5 yellow tiles. One tile is drawn and then replaced. Then a second tile is drawn. What is the probability that the first tile is yellow and the second tile is green? A. 1% B. 3% C. 6% D. 18%

OpenStudy (anonymous):

Total outcome 30 + 15 + 5 = 50 50C1 P(yellow) = 5C1 / 50C1 p(green) = 15C1 / 50C1 Total Prob = P(yellow) + p(green)

OpenStudy (anonymous):

To find % multiply by 100

OpenStudy (campbell_st):

find the probabilities of each colour as a fraction then substitute the specific colour probability into the right hand side and multiply. P(Yellow then Green) = P(Yellow) x P(Green) lastly convert the fraction to a percentage

OpenStudy (anonymous):

huh...Sorry it is P(yellow) x p(green)